OpenStudy (anonymous):

Geometric Sequence: a^1 = 1000 r=2/5 Find explicit expression for A of N, the term number.

OpenStudy (jdoe0001):

\(\Large { a_{\color{brown}{ n}}=a_1\cdot r^{{\color{brown}{ n}}-1}\quad \begin{cases} a_1\to 1^{st}\ term\to 1,000\\ {\color{brown}{ n}}\to n^{th}\ term\to \frac{2}{5} \end{cases} }\)

OpenStudy (anonymous):

Thank you so much! you are whats right in this world :)